OpenStudy (anonymous):

CD is tangent to the circle and segment AE is perpendicular to segment AC. If a radius of the circle measures 12 in., which is the value of DF + AF?

OpenStudy (anonymous):

It uses this picture

OpenStudy (vishal_kothari):

wait for a minute...

OpenStudy (anonymous):

It's multiple choice; these are the answers given: A. 28 + 5(sqrt 3) B. 28 - 4(sqrt 3) C. 24 - 4(sqrt 3) D. 21 + 2(sqrt 3)

OpenStudy (vishal_kothari):

C.

OpenStudy (anonymous):

Can you explain how?

OpenStudy (vishal_kothari):

in which grade you are in?

OpenStudy (anonymous):

8th, but I'm taking Geometry.

OpenStudy (vishal_kothari):

the concept i used is taught in 11th grade...

OpenStudy (anonymous):

Try me

OpenStudy (vishal_kothari):

what?

OpenStudy (anonymous):

As in, let's see if I can understand it.

OpenStudy (vishal_kothari):

in AFE, tan60=AE/AF \[\sqrt{3}\]=12/AF AF=4\[\sqrt{3}\]

OpenStudy (anonymous):

whoops, sorry

OpenStudy (vishal_kothari):

why you are sorry...

OpenStudy (anonymous):

because I accidentally typed a phrase that was incongruous in our conversation?

OpenStudy (anonymous):

oh well

OpenStudy (vishal_kothari):

u in which country...

OpenStudy (anonymous):

America?

OpenStudy (anonymous):

This question was posed to me in my geometry class - are u sure there's no easier way to solve it? I was thinking something to do with 30-60-90 triangles

OpenStudy (vishal_kothari):

i also used that...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well i understand how AF = 4(sqrt 3)

OpenStudy (anonymous):

what im not getting is how to solve for DF

OpenStudy (vishal_kothari):

really...

OpenStudy (vishal_kothari):

u understand how to get AF..

OpenStudy (anonymous):

well, i didn't understand how u explained it, but i did arrive at the same answer - 4(sqrt 3)

OpenStudy (anonymous):

so let's see - if we use 30-60-90 triangles as our basis, then DF = 2CF, correct?

OpenStudy (vishal_kothari):

if i will give you the value of CF.. THAN COULD U FIND DF...

OpenStudy (vishal_kothari):

CF=AC-AF=12-4\[\sqrt{3}\]

OpenStudy (anonymous):

I think I could, but what I'm getting is not 24, like you say it is. That's why I'm going to show you how I would solve it, and I ask that you please correct me if I am wrong. AC = 12, because it is a radius. We have determined that AF = 4(sqrt 3). 12 - 4(sqrt 3) =

OpenStudy (anonymous):

yeah ok good, we agree so far

OpenStudy (anonymous):

so 12 - 4(sqrt 3) = 5.07179677

OpenStudy (anonymous):

Using 30-60-90 Triangles, DF is 2x this, or 10.1435935

OpenStudy (vishal_kothari):

YA..

OpenStudy (vishal_kothari):

ARE U AT FACEBOOK...

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

good, thank you for your time

OpenStudy (vishal_kothari):

WELCOME

OpenStudy (anonymous):

so i am new to this website - how do i thank you in return?

OpenStudy (vishal_kothari):

IT'S 12:25 A.M. HERE..

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

why are you using a computer at 12:30?

OpenStudy (anonymous):

oh well

OpenStudy (anonymous):

so do i thank you by clicking good answer on all of your responses?

OpenStudy (vishal_kothari):

YOUR WISH..

OpenStudy (anonymous):

i told you im new - im not sure what to do now

OpenStudy (vishal_kothari):

WELL IT WAS MY BROTHER'S BIRTHDAY .....SO 12:30..

OpenStudy (vishal_kothari):

GIVE ME A MEDAL AND YOUR LEVEL WILL INCREASE..

OpenStudy (anonymous):

can i give u multiple if i feel like it?

OpenStudy (anonymous):

oh w8, other people have to

OpenStudy (anonymous):

all right, thank you again for all your help

OpenStudy (vishal_kothari):

WELCOME

OpenStudy (vishal_kothari):

ASK MORE QUESTIONS IF YOU HAVE ...

OpenStudy (anonymous):

I'm fine for now, but I definitely will when I have another one.

OpenStudy (vishal_kothari):

OK